1 . Which of the following sets is not closed under addition?
A. odd numbers
B. real numbers
C. whole numbers
D. negative integers
= i got A
2 . Which property is illustrated in the following?
5 + 6 + 7 = 6 + 5 + 7
A. commutative
B. associative
C. reflexive
D. transitive
= i got A
3 . Which property is illustrated in the following?
4 + (3 + x) = (4 + 3) + x
A. commutative
B. reflexive
C. associative
D. transitive
= i got C
4 . Is subtraction of real numbers commutative?
A. Yes
B. No
= i got B
5 . When using a number line to simplify 3 + (–6), you start at which point on the number line?
A. origin
B. –6
C. 3
D. –3
= i got A
6 . Simplify (–4 + 6) – 2.
A. 8
B. 4
C. –4
D. 0
= i got D
7 . Simplify –9 + (–3t) + (–21).
A. –12 + (–3t)
B. –33t
C. –30 + (–3t)
D. 12 + (–3t)
= i got B
8 . Simplify [–7 + 3] + (–4 + (–2)).
A. –2
B. –10
C. –6
D. 2
= i got B
9 . Given x = –3, y = 6, and z = –4, evaluate –z + x – y.
A. –7
B. –13
C. –5
D. –1
= i got C
10 . Simplify x + 5 + [–(x) + (–14)].
A. –9 + 2x
B. 19
C. –9 – x
D. –9
= i got A
Is it all correct ?
Thanks ! :)

Found 2 solutions by rapaljer, jim_thompson5910:
Answer by rapaljer(4671)

I agree with all but 5, 7, and 10.

In #5, I may be wrong, but I would have started at 3. However, it is correct that you have to begin at 0 in order to get to 3. I'm not sure how your particular curriculum views this.

I disagree with #7 and 10. Is this a take home test???

R^2

Dr. Robert J. Rapalje

Answer by jim_thompson5910(35256)

# 1 Correct since odd + odd = even (eg: 3+5=8)

# 2 Correct

# 3 Correct

# 4 Correct (eg: 5-3 does not equal 3-5)

# 5 Incorrect. This question was pretty tricky since you can start at either 3 or -6, but I think the book is looking for you to start with the first given number 3 and then move to the left 6 units (ie subtract 6 from 3). So the answer is C)

# 6 Correct

# 7 Incorrect. You can combine the terms -9 and -21 to get -30, but you CANNOT combine -30 and -3t to get -33t since they are NOT like terms. The only thing you can do is to leave them as they are and say that the answer is -30+(-3t) which means that the answer is C)

# 8 Correct.

# 9 Correct.

# 10 Incorrect. Note how x + (-x) is the same as x - x which subtracts to 0x or just 0. This means that the 'x' terms just cancel out and go away. We're then left with 5 + (-14) = -9 which gives the answer D) -9